# Probing Non-Thermal Processes in Stellar Flares on AU Mic

### v1

@@ 4.1 The Orrall-Zirker Effect
The Ly emission line was masked in the observations to avoid detector saturation from 1206 - 1220 Å. Even with this mask, the spectral range is sufficient to search for evidence of this effect in the HST/COS data set. Should a non-thermal proton beam be present during these flares, we would expect to see the red wing emission enhancement during the early impulsive phase of a flare (Woodgate et al., 1992), when much of the atmosphere among the flaring areas is neutral. Thus, we calculated spectroscopic light curves of the red wing (1220 - 1233.37 Å) and blue wing (1197.97 - 1206 Å) of the Ly emission line to detect the OZ effect. Within the blue wing, we masked S III ( = 1200.97, 1206.53, 1207.51 Å), as identified using the HST/STIS spectrum of AU Mic presented in Rockcliffe et al. (2023a). This left a sufficient baseline to determine if the blue wing responded asymmetrically to the red wing in the observed flares. This analysis was repeated for six of the highest-energy flares in our data set: Flares B, J, K, M, R, and T .
@@ 4.2 Observed Carbon Blue Enhancement
Flare B is a large triple-peaked flare event and the longest-lived event observed in the FUV. We present the blue wing, core, and red wing light curves of C II and C III in Figures 4 and 5, respectively. Both figures show that the blue wings of C II and C III appear to respond mainly to the first peak in Flare B, which is further evidence that the second peak in Flare B is a second flare due to its different spectral responses. In the second peak, the blue enhancement of the C III and C II spectral lines is not as pronounced as in the first peak. The line response in the second peak seems to be similar to other lines during the flare. In fact, we see in C II and C III that the second flare in the Flare B event has a stronger response in the red wing than the first flare. This highlights the unique emission line response for sequential flares. This response was analyzed in other C emission lines, but these emission lines lacked the signal-to-noise needed to make a clear observation of the spectral response of the blue and red wings throughout Flare B.
@@ 4.3 Modeling Chromospheric Evaporation
Chromospheric evaporation occurs when non-thermal electrons stream along magnetic field lines and impact the chromosphere. Once these non-thermal electrons arrive at the chromosphere, they deposit their energy into the plasma, heating the plasma and increasing pressure. The heated plasma is accelerated upwards, resulting in chromospheric evaporation Fisher et al. (1985). This phenomenon is observed in the Sun. Observational evidence of chromospheric evaporation includes blue-shifted spectral lines during the impulsive phase of the flare (Milligan & Dennis, 2009; Antonucci et al., 1984; Shibata et al., 1995). If the blue-enhancement seen in C II and C III is due to chromospheric evaporation, this would imply that the gas emitted at K is being accelerated upwards through the flare loop. The blueshift would be seen at the foot points of the soft X-ray (SXR) loop depending on its temperature. Here, we aim to determine if models of solar-like chromospheric evaporation can be used to explain the observed C II blue enhancement.
@@
We used a 1D radiative-hydrodynamic M-dwarf stellar flare model from the grid described in Kowalski et al. (2024) to model the spectrum of C II during Flare B. We generated a flare atmosphere model with the following assumptions: (i) peak injected electron beam energy flux density of [erg cm-2 s-1], (ii) a low energy cutoff of 37 keV, and (iii) a power law index of = 2.5. We chose this model because it is able to reproduce M dwarf optical hydrogen lines well when combined with a model for the continuum (Kowalski, 2022; Kowalski et al., 2025). The chosen model also produces upflows of several hundred km s-1 that could inform our understanding of mass motions during the flare events. This model is a higher energy extension of the F11 solar flare model in Allred et al. (2005), which produces prominent blueshifts in cool temperature ( K) flare lines.
@@
We modeled individual time steps between the impulsive and decay phases of Flare B using the RH radiative transfer code from Uitenbroek (2001). We evaluated the model for Flare B between seconds at s time-intervals. Note that the model timescale is different than the timescale of Flare B, which lasts seconds (2). The models in the grid represent a superposition of short bursts; nonetheless analysis of individual time steps is necessary to understand how the atmosphere responds in a spatially unresolved observation. To synthesize the C II emission line, we use the model atom from (Rathore et al., 2015). A handful of time steps required modification to the number of depth points () included in the model. Initially, the model had . We had to reduce to 180 in most cases, excluding heights in the upper corona, to achieve convergence. As the populations of C II are very small in these hot regions, truncating the atmosphere in this way does not affect our results. Additionally, time steps from s required . We concluded this would be an insufficient height representation for an accurate stellar flare model. Therefore, we do not include models at s for C II in our analysis. All models are calculated at 5 emergent rays ( 18.19∘, 38.74∘, 60.00∘, 76.70∘, and 87.13∘), and are summed using Gaussian quadrature to give an emergent radiative flux spectrum.
@@
The model comparisons are shown in Figure 6. While every model has limitations, we can draw several conclusions concerning our observation of the response of the C II emission lines during Flare B from the Kowalski et al. (2024) model. First, we find that the indirect emerging ray ( = 87.13∘) contributes to line broadening, while the direct emerging ray ( = 18.19∘) contributes to the spectral line strength. However, none of the rays or the flux account for the discrepancy between the blue and red enhancements of the C II spectral line; there is no significant redshift or blueshift in the model for C II.
@@ 5 Discussion
In Ikuta & Shibata (2024), a one-dimensional free-fall model and a pseudo-2D MHD model were used to infer the direction of the ejection, changes in height, and the possible development of a magnetic loop during a flare on EK Draconis. A significant redshift of the H spectral line was predicted; this magnitude of redshift is not present in the C II or C III spectral lines during Flare B. We speculate that the blue enhancement observed here for C III and C II could be explained by emission coming from the expansion of the legs of the filament. This expansion would produce an initial blueshift, followed by a redshift as the filament expands. The phenomenon was reported for the H line during a superflare on EK Draconis, in Ikuta & Shibata (2024) the H line variance was modeled during a flare, and it was found that filament eruptions could explain the variance observed. The blue enhancement caused by the filament eruptions has a particular velocity profile starting at velocities of a few tens of km/s up to hundreds of km/s back to zero, and then there is a redshift later on; this process is illustrated in Figure 10 of Ikuta & Shibata (2024). In Ikuta & Shibata (2024), there was an initial acceleration of over 500 km/s for the H spectral line, decelerating to -200 km/s. During Flare B, there was an initial similar blue acceleration of C II and C III, in the range of a few hundred km/s.
@@
Current abilities to model broadening of Mg II for solar flares fall drastically short (Zhu et al., 2019). Our research highlights that a similar problem potentially exists for C II. Zhu et al. (2019) introduced a Stark width factor of 30. We explore the role of this for C II, shown in Figure 6. By multiplying by a factor of 30 for the C II resonance lines, the model is better able to describe the broadening throughout the flare; this suggests that an unknown source of atomic line broadening may also explain the C II widths in stellar flares. The enhanced symmetric broadening extends far into the wings (Figure 6, s). Determining the physical origin of the missing amounts of the broadening would thus improve the interpretation of asymmetric wings due to mass motions. In future work, we plan to explore the role of this parameter in more detail while including the C III lines (which are not included in the Rathore et al. 2015 model atom) in the calculations.
5.1 FUV Flare Fitting
@@
The flare energies and equivalent durations presented here were derived from the light curve. Other work have reported these values derived from a best-fit model. We attempted to model these FUV flares with the open-source code Llamaradas Estelares (Tovar Mendoza et al., 2022), an analytical flare model that was optimized for optical flares. We present an example fit to Flare R in Figure 7. The Llamaradas Estelares flare model was unable to capture the decay phases of several flares, including Flare R, accurately. In particular, the model was unable to account for the extended tail of Flare R, which could be due to subsequent sympathetic flares or persistent emission lines during the decay phase. We emphasize exercising caution in future work when using this model to calculate flare energy and duration, as it has not consistently provided an accurate fit in this set of flares. A more flexible model is needed for modeling FUV flares.
@@ 6 Conclusions and Future Work
We present the first comparison of the FUV and optical flare frequency distributions (FFDs) of AU Mic (Section 3.2). The differences in the FFDs suggest that the average energy budget of the flares dominates in the FUV over the optical, which is consistent with our understanding of flare temperatures.
We find no evidence of the OZ effect in the high-energy flares analyzed (Section 4); Woodgate et al. (1992) remains the only tentative detection of the OZ effect on another star to-date. Our non-detections provide insight into the frequency of proton beams in M-dwarf flares and can be used to enhance our models of such events, which currently lack energy transport from non-thermal protons.

### v2

@@ 4.1 The Orrall-Zirker Effect
The Ly emission line was masked in the observations to avoid detector saturation from 1206 - 1220 Å. Even with this mask, the spectral range is sufficient to search for evidence of this effect in the HST/COS data set. Should a non-thermal proton beam be present during these flares, we would expect to see the red wing emission enhancement during the early impulsive phase of a flare (Woodgate et al., 1992), when much of the atmosphere among the flaring areas is neutral. Thus, we calculated spectroscopic light curves of the red wing (1220 - 1233.37 Å) and blue wing (1197.97 - 1206 Å) of the Ly emission line to detect the OZ effect. Within the blue wing, we masked S III ( = 1200.97 Å) and Si III ( = 1206.53, 1207.51 Å), as identified using the HST/STIS spectrum of AU Mic presented in Rockcliffe et al. (2023a). This left a sufficient baseline to determine if the blue wing responded asymmetrically to the red wing in the observed flares. This analysis was repeated for six of the highest-energy flares in our data set: Flares B, J, K, M, R, and T. From the flare properties calculated in Section 3, we can determine that these 6 high-energy flares have Si III energy the Si III energy from the flare identified in Woodgate et al. (1992), where the OZ effect was first tentatively detected.
@@ 4.2 Observed Carbon Blue Enhancement
Flare B is a large triple-peaked flare event and the longest-lived event observed in the FUV. We present the blue wing, core, and red wing light curves of C II and C III in Figures 3 and 4, respectively. Both figures show that the blue wings of C II and C III appear to respond mainly to the first peak in Flare B, which is further evidence that the second peak in Flare B is a second flare due to its different spectral responses. In the second peak, the blue enhancement of the C III and C II spectral lines is not as pronounced as in the first peak. The line response in the second peak seems to be similar to other lines during the flare. In fact, we see in C II and C III that the second flare in the Flare B event has a stronger response in the red wing than the first flare. This highlights the unique emission line response for sequential flares.
We searched for evidence of blue wing enhancements across all other emission lines in our observations. For emission lines with sufficient signal-to-noise (e.g., N V, Ly wings), we were unable to detect any asymmetrical responses between the blue and red wings of those lines. All other emission lines lacked the signal-to-noise to sufficiently model the line profile.
@@ 4.3 Modeling Chromospheric Evaporation
If the blue-enhancement seen in C II and C III is due to chromospheric evaporation, this would imply that the gas emitted at K is being accelerated upwards through the flare loop. The blueshift would be seen at the foot points of the soft X-ray (SXR) loop depending on its temperature. Here, we aim to determine if models of solar-like chromospheric evaporation can be used to explain the observed C II blue enhancement.
@@
We used a 1D radiative-hydrodynamic M-dwarf stellar flare model from the grid described in Kowalski et al. (2024) to model the spectrum of C II during Flare B. We generated a flare atmosphere model with the following assumptions: (i) peak injected electron beam energy flux density of [erg cm-2 s-1], (ii) a low energy cutoff of 37 keV, and (iii) a power law index of = 2.5. We chose this model because it is able to reproduce M dwarf optical hydrogen lines well when combined with a model for the continuum (Kowalski, 2022; Kowalski et al., 2025). The chosen model also produces upflows of several hundred km s-1 that could inform our understanding of mass motions during the flare events. This model is a higher energy extension of the F11 solar flare model in Allred et al. (2005), which produces prominent blueshifts in cool temperature (T K) flare lines.
@@
We modeled individual time steps between the impulsive and decay phases of Flare B using the RH radiative transfer code from Uitenbroek (2001). We evaluated the model for Flare B between t seconds at t second time-intervals. Note that the model timescale is different than the timescale of Flare B, which lasts seconds (1). The models in the grid represent a superposition of short bursts; nonetheless analysis of individual time steps is necessary to understand how the atmosphere responds in a spatially unresolved observation. To synthesize the C II emission line, we use the model atom from (Rathore et al., 2015). A handful of time steps required modification to the number of depth points () included in the model. Initially, the model had . We had to reduce to 180 in most cases, excluding heights in the upper corona, to achieve convergence. As the populations of C II are very small in these hot regions, truncating the atmosphere in this way does not affect our results. Additionally, time steps from s required . We concluded this would be an insufficient height representation for an accurate stellar flare model. Therefore, we do not include models at s for C II in our analysis. All models are calculated at 5 emergent rays ( 18.19∘, 38.74∘, 60.00∘, 76.70∘, and 87.13∘), and are summed using Gaussian quadrature to give an emergent radiative flux spectrum.
@@
The model comparisons are shown in Figure 5. While every model has limitations, we can draw several conclusions concerning our observation of the response of the C II emission lines during Flare B from the Kowalski et al. (2024) model. First, we find that the indirect emerging ray ( = 87.13∘) contributes to line broadening, while the direct emerging ray ( = 18.19∘) contributes to the spectral line strength. However, none of the rays or the flux account for the discrepancy between the blue and red enhancements of the C II spectral line; there is no significant redshift or blueshift in the model for C II. This chromospheric evaporation model does not explain the observed blue enhancement of the C II spectral line, which could be indicative of a lack of chromospheric evaporation during this flare or a deficiency in current models.
@@ 5 Discussion
In Ikuta & Shibata (2024), a one-dimensional free-fall model and a pseudo-2D MHD model were used to infer the direction of the ejection, changes in height, and the possible development of a magnetic loop during a flare on EK Draconis. A significant redshift of the H spectral line was predicted; this magnitude of redshift is not present in the C II or C III spectral lines during Flare B. We speculate that the blue enhancement observed here for C III and C II could be explained by emission coming from the expansion of the legs of the filament. Solar IRIS observations reveal that C II and Mg II exhibit blue shifts attributed to chromospheric evaporation Li et al. (2019). Additional analyses of IRIS observations of these spectral lines may provide further insights into spectral tracers of filament eruptions (e.g., Panos et al. (2018)). This expansion would produce an initial blueshift, followed by a redshift as the filament expands. The phenomenon was reported for the H line during a superflare on EK Draconis, in Ikuta & Shibata (2024) the H line variance was modeled during a flare, and it was found that filament eruptions could explain the variance observed. The blue enhancement caused by the filament eruptions has a particular velocity profile starting at velocities of a few tens of km/s up to hundreds of km/s back to zero, and then there is a redshift later on; this process is illustrated in Figure 10 of Ikuta & Shibata (2024). In Ikuta & Shibata (2024), there was an initial acceleration of over 500 km/s for the H spectral line, decelerating to -200 km/s. During Flare B, there was an initial similar blue acceleration of C II and C III, in the range of a few hundred km/s.
@@
Similar models to those used here have been employed for solar flares, and similar line broadening modeling deficiencies have been identified in other spectral lines, such as the Mg II resonance lines in the near UV (Zhu et al., 2019; Hawley et al., 2007). Current abilities to model broadening of Mg II for solar flares fall drastically short (Zhu et al., 2019). Our research highlights that a similar problem potentially exists for C II. Zhu et al. (2019) introduced a Stark width factor of 30. We explore the role of this for C II, shown in Figure 5. By multiplying by a factor of 30 for the C II resonance lines, the model is better able to describe the broadening throughout the flare; this suggests that an unknown source of atomic line broadening may also explain the C II widths in stellar flares. The enhanced symmetric broadening extends far into the wings (Figure 5, t s). Determining the physical origin of the missing amounts of the broadening would thus improve the interpretation of asymmetric wings due to mass motions. In future work, we plan to explore the role of this parameter in more detail while including the C III lines (which are not included in the Rathore et al. 2015 model atom) in the calculations.
@@
The flare energies and equivalent durations presented here were derived from the light curve. Other work have reported these values derived from a best-fit model. We attempted to model these FUV flares with the open-source code Llamaradas Estelares (Tovar Mendoza et al., 2022), an analytical flare model that was optimized for optical flares. We present an example fit to Flare R in Figure 6. The Llamaradas Estelares flare model was unable to capture the decay phases of several flares, including Flare R, accurately. In particular, the model was unable to account for the extended tail of Flare R, which could be due to subsequent sympathetic flares or persistent emission lines during the decay phase. We emphasize exercising caution in future work when using this model to calculate flare energy (E) and duration, as it has not consistently provided an accurate fit in this set of flares. A more flexible model is needed for modeling FUV flares.
@@ 6 Conclusions and Future Work
We find no evidence of the OZ effect in the high-energy flares analyzed (Section 4.1); Woodgate et al. (1992) remains the only tentative detection of the OZ effect on another star to-date. Our non-detections provide insight into the frequency of proton beams in M-dwarf flares and can be used to enhance our models of such events, which currently lack energy transport from non-thermal protons.
@@
The HST data presented in this article was obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/bh63-4j34 (catalog doi: 10.17909/bh63-4j34).